# Mut4All: Fuzzing Compilers via LLM-Synthesized Mutators Learned from Bug Reports

### v1

@@ 4.5.1. Overall Effectiveness and Cost
For Rust, Mut4All generated 64% valid mutators (319 out of 500). Among invalid ones, 110 were filtered out for producing no observable mutations on the test suite, and 71 failed to compile after 10 rounds of refinement. Notably, 397 of the Rust mutators required at least one round of refinement, with an average of 3.51 refinement iterations per mutator. On average, generating one valid Rust mutator consumed 14,969 input tokens and 3,694 output tokens, corresponding to \$0.074 each based on the GPT-4o pricing in 2025.
For C++, Mut4All synthesized 403 valid mutators (81%), while 9 were ineffective and 88 failed after 10 refinement attempts. Among these, 264 (52.8%) required refinement, with an average of 2.48 iterations. The generation of each valid C++ mutator consumed an average of 22,276 input tokens and 2,500 output tokens, amounting to a cost of approximately \$0.080.
Overall, Mut4All achieved a higher success rate on C++ than on Rust, primarily due to C++’s more concise and high-level AST transformation APIs. In contrast, Rust’s verbose, low-level visitor-based traversal increases the likelihood of failure. Importantly, since the entire synthesis process is fully automated and requires no human intervention, the cost of approximately \$0.08 per valid mutator is highly affordable and practical for large-scale use.
@@ 4.5.2. Analysis of Valid Mutators
We manually analyzed 65 Rust and 68 C++ mutators, a statistically representative number with a 90% confidence level and a 10% error margin. Two authors manually review each mutator and examined the alignment between code logic and the corresponding specifications.
The analysis showed that only 35% of the Rust mutators and 28% of the C++ mutators faithfully implemented the intended mutations. Interestingly, among the mutators that triggered confirmed bug discoveries, approximately 65% were fully aligned with their specifications. This suggests that even partially incorrect mutators can uncover bugs! Nonetheless, higher implementation fidelity is generally associated with improved bug-finding ability, underscoring the importance of enhancing mutator correctness in future work.
To further understand the the errors in the mutators, which guide the direction to improve our method, we categorized the implementation errors into three types: (1) Missing changes (19%): the specification involves changes at multiple locations, but some were omitted; (2) Incorrect mutation (45%): the code transformation was implemented incorrectly; (3) Incorrect matching (36%): the mutator failed to match the correct AST nodes.
@@ 4.5.3. Analysis of Invalid Mutators
Using the same sampling parameters (90% confidence level, 10% error margin), we analyzed 45 Rust and 43 C++ invalid mutators. Via open coding, we identified 5 primary causes of failure: (1) Use of deprecated APIs (40%), (2) Incorrect parameters (39%), (3) Syntax parsing errors (10%), (4) Operand type mismatch (7%), and (5) Use of undeclared names (5%).
The most common failure in Rust was Incorrect parameters (62%), while in C++ it was the Use of deprecated APIs (65%). These results indicate that synthesizing structurally correct and type-safe AST manipulations remains challenging for advanced LLMs, particularly in languages with verbose or rapidly evolving APIs like Rust.
Answer to RQ3: Mut4All automatically synthesizes valid and effective mutators with low cost (about \$0.08 each) without human intervention. Improving the correctness of LLM-synthesized mutators remains promising for future work.
@@ 5.1. Limitation and Implication
Despite these advantages, our evaluation also reveals several key limitations. First, only about 30% of the valid mutators fully adhere to their intended specifications. Among those that triggered confirmed bugs, about 65% were correctly implemented. This suggests that while imperfect mutators may still uncover bugs by chance, higher implementation accuracy generally correlates with better bug-finding effectiveness. Improving the semantic fidelity of generated mutators remains a critical direction for future work. Moreover, many invalid mutators fail due to deprecated APIs, likely inherited from legacy training data. To enhance correctness, future work could explore symbolic reasoning or reinforcement learning and fine-tune models on modern code.
@@ 5.2. Threats to Validity
Internal Validity.  Our pipeline depends on LLM output quality and prompt design. Although we employ validation and refinement to ensure correctness, LLM-generated mutators may still misbehave in edge cases. Manual analysis may also introduce bias, which we mitigate using statistically sound sampling (90% confidence, 10% margin of error).
External Validity.  We evaluate on Rust and C++, using four widely adopted compilers. While representative, results may not generalize to other languages or domains. Baselines span both traditional and LLM-based fuzzers, but future tools may differ in design.
Construct Validity.  We primarily use crash count and line coverage as evaluation metrics. While standard in compiler fuzzing, these do not capture semantic correctness or non-crashing logic bugs. Additionally, fuzzers are executed under timeout constraints, which may introduce variance; we mitigate this by averaging results across multiple independent runs, i.e., 5 runs for each fuzzer.
@@ 6.1. Compiler Fuzzers
Generation-based fuzzers synthesize programs from scratch using grammar rules or probabilistic models. Tools such as CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023) exemplify this paradigm by generating semantically valid C/C++ code. Recent extensions have expanded to other languages, including Java (Chen et al., 2023; Zhao et al., 2022), JavaScript (Ye et al., 2023), Rust (Sharma et al., 2023; Wang and Jung, 2024), MLIR (Wang et al., 2023a), deep learning compilers (Wang et al., 2023b; Liu et al., 2023a; Ren et al., 2025), SQL (Hua et al., 2023), and Kotlin (Georgescu et al., 2024). LLM-based generators (Yang et al., 2024a; Gao et al., 2025) further automate program synthesis but struggle with intricate language constructs (e.g., generics, trait bounds), which are difficult to generate from scratch while preserving semantic correctness. In contrast, Mut4All mutates real-world programs—retaining structural complexity and ensuring semantic validity—thereby enabling deeper and more effective exploration of compiler internals.
Mutation-based fuzzers generate program variants by applying transformations to seed programs. Representative approaches include EMI-based fuzzers (Le et al., 2014, 2015; Gao et al., 2024b) and differential fuzzers targeting C/C++ (Even-Mendoza et al., 2023; Jiang et al., 2021; Liu et al., 2023b), JIT languages (Chen et al., 2016b, 2019; Schumi and Sun, 2021; Wu et al., 2023; Park et al., 2023; Gao et al., 2024a), deep learning compilers (Li et al., 2023; Chen et al., 2025b), MLIR (Suo et al., 2024), Rust (Yang et al., 2024b), and various domain-specific languages (Tu et al., 2022; Bendrissou et al., 2024). These methods typically rely on manually designed mutators, requiring significant domain expertise and engineering effort. Among them, MetaMut (Ou et al., 2024) represents the first attempt to synthesize mutators using LLMs, composing transformations from predefined AST-level operations. However, its search space is constrained by these predefined operations, and over half of the generated mutators require manual repair to function correctly. In contrast, Mut4All automates both the invention and implementation of mutators through a multi-agent LLM framework guided by real-world bug reports crawled from open-source compiler communities. By grounding mutation design and synthesis in authentic bug-prone features and enabling collaboration among specialized LLM agents, Mut4All explores a broader mutator design space with greater creativity and incorporates automatic validation and refinement mechanisms to ensure correctness, significantly reducing human effort while improving performance.
@@ 6.2. LLM-Based Testing
Recent advances have demonstrated the potential of LLMs in automated software testing (Wang et al., 2024; Chen et al., 2025a). Many works focus on generating unit tests, including test functions (Schäfer et al., 2023; Chen et al., 2024a; Yang et al., 2024c; Yuan et al., 2024; Shang et al., 2025), input values (Chen et al., 2024b), and assertion oracles (Hossain et al., 2025; Molina et al., 2025), often guided by code structure, comments, or coverage feedback. Bug reports are used to enhance test generation (Ouédraogo et al., 2025). These approaches primarily operate at the unit-test level, aiming to validate method-level behavior, and their outputs are usually test code written in the same language as the program under test. In contrast, our work targets compiler testing, focusing on mutation operators that systematically mutate code to explore deep compiler behavior.
@@ 7. Conclusion
We present Mut4All, a fully automated, language-agnostic framework that synthesizes mutation operators using LLMs guided by historical bug reports. Mut4All uses three LLM agents for mutator invention, implementation, and refinement. It generates 319 and 403 valid mutators for Rust and C++, respectively, which uncovered 96 compiler bugs, 54 of which were previously unknown. The low cost demonstrates the promise of LLM-enhanced pipelines as a practical alternative to manual mutator development. In the future, we plan to reduce invalid mutators and enhance the correctness of mutator implementation by symbolic reasoning or reinforcement learning.

### v2

@@ 4.6. RQ4: Mutator Synthesis Effectiveness
We classify mutators as valid or invalid based on whether they can be successfully compiled and correctly applied to all programs in our test suite. For Rust, Mut4All generated 64% valid mutators (319 out of 500). Among the invalid ones, 110 were filtered out for producing no observable mutations on the test suite, and 71 failed to compile after 10 rounds of refinement. Notably, 397 of the Rust mutators required at least one round of refinement, with an average of 3.51 refinement iterations per mutator. For C++, Mut4All synthesized 403 valid mutators (81%), while 9 had no effect agasint the test suite and 88 failed after 10 rounds of refinement attempts. Among these, 264 (52.8%) required at least one attempt of refinement, with an average of 2.48 iterations.
We further manually analyze the characteristics of both valid and invalid mutators to better understand their behaviors.
4.6.1. Labeling Process
To better understand the behavior and quality of synthesized mutators, we conduct a manual inspection over a statistically representative subset of both valid and invalid mutators. Specifically, to ensure the robustness of our analysis, we sample mutators at a 95% confidence level with a 5% margin of error.
We apply a structured card sorting methodology to both categories (Yang et al., 2024d; Wang et al., 2025a). (1) For valid mutators, we verify whether the transformation fully conforms to the intended mutation specification synthesized in the earlier pipeline stage. If the output does not completely fulfill the semantic intent, we classify it as a specification mismatch and further group the failure causes. (2) For invalid mutators, we similarly categorize the failure reasons. Each mutator is independently assessed by at least two annotators, with conflicts resolved through consensus.
@@ 4.6.2. Analysis of Valid Mutators
We sampled 175 Rust and 197 C++ valid mutators, ensuring the 95% confidence level and 5% error margin. Two authors manually reviewed each mutator and examined the alignment between code logic and the corresponding specifications. The analysis showed that 44% of the Rust mutators and 37% of the C++ mutators faithfully implemented the intended mutation specifications. Interestingly, among the mutators that triggered confirmed bug discoveries, approximately 65% were fully aligned with their specifications. This suggests that even partially incorrect mutators can uncover bugs! Nonetheless, higher implementation fidelity is generally associated with improved bug-finding ability, underscoring the importance of enhancing mutator correctness in future work.
To further understand the errors in the mutators, which guide the direction to improve our method, we categorized the implementation errors into three types.
(1) Incorrect AST Node Matching: the mutator failed to match the correct AST nodes. For example, the LLM-synthesized mutators usually mistakenly use the variable names that are used in the specification examples, rather than the AST node type of identifier, to locate the target node. This type accounts for 11% and 22% of the valid mutators for Rust and C++, respectively.
(2) Incorrect Mutation: the mutator incorrectly implements the intended transformation described in the specification. For instance, the specification may require deleting only the right-hand side expression of an assignment statement. However, the LLM-synthesized mutator mistakenly removes all tokens following the assignment operator, including the semicolon at the end of the line, resulting in unintended behavior. This type accounts for 31% and 31% of the valid mutators for Rust and C++, respectively.
(3) Missing Steps of Specification: the specification requires multiple coordinated edits, but the mutator only performs a subset of them. For example, a specification may instruct the removal of a parameter from a function signature. However, the LLM‑synthesized mutator may only modify the function signature while failing to remove the parameter’s remaining references, resulting in semantic errors. This type accounts for 14% and 10% of the valid mutators for Rust and C++, respectively.
@@ 4.6.3. Analysis of Invalid Mutators
Using the same sampling setups, we analyzed 61 Rust and 72 C++ invalid mutators. We identified 5 primary causes of failure:
  1. (1)
    Use of Deprecated APIs, the mutator fails because it invokes outdated AST manipulation APIs. These APIs have changed with evolution, but the LLM may have learned from legacy training data. Therefore, it generates transformations using API calls that no longer exist or behave differently, leading to errors. This type accounts for 15% and 67% of the invalid mutators for Rust and C++, respectively.
  2. (2)
    Incorrect Parameters, the mutator fails due to the misuse of APIs with invalid arguments. This type accounts for 51% and 17% of the invalid mutators for Rust and C++, respectively.
  3. (3)
    Syntax Errors, the mutator fails due to syntactic errors within its implementation. This type accounts for 18% and 8% of the invalid mutators for Rust and C++, respectively.
  4. (4)
    Operand Type Mismatch, the mutator fails due to the misuse of the operands of incorrect types. This type accounts for 15% and 1% of the invalid mutators for Rust and C++, respectively.
  5. (5)
    Use of Undeclared Names, the mutator fails because it uses variables or functions without declarations. This type accounts for 1% and 7% of the invalid mutators for Rust and C++, respectively.
The most common failure in Rust was Incorrect parameters (51%), while in C++ it was the Use of deprecated APIs (67%). These results indicate that synthesizing structurally correct and type-safe AST manipulations remains challenging for advanced LLMs, particularly in languages with verbose or rapidly evolving APIs like Rust.
Answer to RQ4: For the valid mutators, 44% of the Rust mutators and 37% the C++ mutators faithfully implemented the intended mutation specifications. Incorrect Mutation is the most common cause of the incorrect implementation. For the invalid mutators, Incorrect Parameters and Use of Deprecated APIs are the most common causes for compilation errors in Rust and C++, respectively.
@@ 5. Discussion
In this section, we first discuss the limitations of our study and the implications for future work, then discuss the threats to validity.
@@ 5.1. Limitation and Implication
Despite these advantages, our evaluation also reveals several key limitations. First, only about 40% of the valid mutators fully adhere to their intended specifications. Among those that triggered confirmed bugs, about 65% were correctly implemented. This suggests that while imperfect mutators may still uncover bugs by chance, higher implementation accuracy generally correlates with better bug-finding effectiveness. Improving the semantic fidelity of generated mutators remains a critical direction for future work. Additionally, our analysis of invalid mutators shows that many failures stem from the use of deprecated APIs—likely a result of legacy patterns present in the LLM’s training corpus. Future work should explore techniques to proactively detect and avoid such outdated constructs during generation, such as compiler warning filtering, specification-grounded constraints, or fine-tuning on modern codebases.
@@ 5.2. Threats to Validity
5.2.1. Internal Validity
Our pipeline depends on LLM output quality and prompt design. Although we employ validation and refinement to ensure correctness, LLM-generated mutators may still misbehave in edge cases. Manual analysis may also introduce bias, which we mitigate using statistically sound sampling (95% confidence, 5% margin of error).
5.2.2. External Validity
We evaluate our approach on Rust and C++, two widely-used system-level programming languages, and using four widely adopted, production-level compilers. While representative, results may not generalize to other languages or domains. Baselines span both traditional and LLM-based fuzzers, but future tools may differ in design.
5.2.3. Construct Validity
We primarily use crash count and line coverage as evaluation metrics. While standard in compiler fuzzing, these do not capture semantic correctness or non-crashing logic bugs. Additionally, fuzzers are executed under timeout constraints, which may introduce variance; we mitigate this by averaging results across multiple independent runs, i.e., 5 runs for each fuzzer.
@@ 6. Related Work
In this section, we discuss the related work of our study.
@@ 6.1. Compiler Fuzzers
6.1.1. Generation-based fuzzers
Generation-based fuzzers synthesize programs from scratch, typically using predefined grammar rules or learned probabilistic models. Tools like CSmith (Yang et al., 2011) and YARPGen (Livinskii et al., 2020, 2023) exemplify this category, generating semantically valid programs to explore diverse compiler behaviors. More recent work has extended this paradigm into other languages, including Java (Chen et al., 2023; Zhao et al., 2022), JavaScript (Ye et al., 2023), Rust (Sharma et al., 2023; Wang and Jung, 2024), MLIR (Wang et al., 2023a), deep-learning compilers (Wang et al., 2023b; Liu et al., 2023b; Ren et al., 2025), SQL executors (Hua et al., 2023), and Kotlin (Georgescu et al., 2024; Chaliasos et al., 2022). Researchers also proposed auxiliary approaches to boost generation-based fuzzers (Even-Mendoza et al., 2020, 2022; Wu et al., 2023a; Li et al., 2024). The most recent generation-based approaches leverage LLMs for generating code (Yang et al., 2024a; Gao et al., 2025; Ni and Li, 2025). However, a key limitation of generation-based methods is their inability to synthesize complex programs involving complex language features (e.g., generics, trait bounds, etc.). These constructs are often too fragile to be generated from scratch, especially with LLMs’ limited control over semantic consistency. In contrast, Mut4All operates by mutating real-world seed programs and thus naturally preserves complex syntactic and semantic structures.
6.1.2. Mutation-based fuzzers
Mutation-based fuzzers generate program variants by applying transformations to seed programs. Representative approaches include EMI-based fuzzers (Le et al., 2014, 2015; Gao et al., 2024a; Theodoridis et al., 2022; Sun et al., 2016) and differential fuzzers targeting C/C++ (Even-Mendoza et al., 2023; Jiang et al., 2021; Liu et al., 2023a; Wang et al., 2025b; Wu et al., 2025), JIT languages (Chen et al., 2016b, 2019; Schumi and Sun, 2021; Wu et al., 2023b; Park et al., 2023; Zang et al., 2023; Xie et al., 2024; Feng et al., 2025), deep learning compilers (Li et al., 2023; Chen et al., 2025b), MLIR (Suo et al., 2024), Rust (Yang et al., 2024c), and various domain-specific languages (Tu et al., 2022; Bendrissou et al., 2024). These methods typically rely on manually designed mutators, requiring significant domain expertise and engineering effort. Among them, MetaMut (Ou et al., 2024) represents the first attempt to synthesize mutators using LLMs, composing transformations from predefined AST-level operations. However, its search space is constrained by these predefined operations, and over half of the generated mutators require manual repair to function correctly.
In contrast, Mut4All automates both the invention and implementation of mutators through a multi-agent LLM framework guided by real-world bug reports crawled from open-source compiler communities. By grounding mutation design and synthesis in authentic bug-prone features and enabling collaboration among specialized LLM agents, Mut4All explores a broader mutator design space with greater creativity and incorporates automatic validation and refinement mechanisms to ensure correctness, significantly reducing human effort while improving performance.
@@ 6.2. LLM-Based Testing
Recent advances have demonstrated the potential of LLMs in automated software testing (Wang et al., 2024; Chen et al., 2025a). Many works focus on generating unit tests, including test functions (Schäfer et al., 2023; Chen et al., 2024b; Yang et al., 2024b; Yuan et al., 2024; Shang et al., 2025), input values (Chen et al., 2024a), and assertion oracles (Hossain et al., 2025; Molina et al., 2025), often guided by code structure, comments, or coverage feedback. Bug reports are used to enhance test generation (Ouédraogo et al., 2025). These approaches primarily operate at the unit-test level, aiming to validate method-level behavior, and their outputs are usually test code written in the same language as the program under test. In contrast, our work targets compiler testing, focusing on mutation operators that systematically mutate code to explore deep compiler behavior.
@@ 7. Conclusion
We present Mut4All, a fully automated, cross-language framework that synthesizes mutation operators using LLMs guided by historical bug reports. Mut4All involves three LLM agents, namely Mutator Invention agent, Mutator Implementation Synthesis agent, and Mutator Refinement agent. It generates 319 and 403 valid mutators for Rust and C++, respectively. Moreover, we expand the input seed programs with our adaptive seed enhancement approach. In our bug-hunting experiment, Mut4All uncovered 96 compiler bugs, 58 of which were previously unknown and missed by other compiler fuzzers. The low cost demonstrates the promise of LLM-enhanced pipelines as a practical alternative to manual mutator development. In the future, we plan to reduce invalid mutators and enhance the correctness of mutator implementation by symbolic reasoning or reinforcement learning. Our artifact is publicly available at https://github.com/sososopy/Mut4All-Artifacts.